OpenStudy (anonymous):

If f(x) = 8 arctan(7 e^x), find f'( x ). Could sme please tell me the procedure and answer?

OpenStudy (zzr0ck3r):

do you know what the chain rule is?

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

if\[f(x) = g(h(x))\\f'(x) = g'(h(x))*h'(x)\]

OpenStudy (zzr0ck3r):

what is the derivative of arctan(x)?

OpenStudy (anonymous):

1/(1+x^2)

OpenStudy (zzr0ck3r):

ok so you have \[f(x) = 8\arctan(g(x))\] where g(x) = 7e^x so \[f'(x) = 8*\frac{1}{1+g(x)^2}*g'(x) \] replace g(x) with 7e^x \[f'(x)=\frac{8}{1+(7e^x)^2}*(7e^x)'=\frac{8}{1+49e^{2x}}*7e^x=\frac{56e^x}{1+49e^{2x}}\]

OpenStudy (anonymous):

thank you for your help! I really appreciate it . I just forgot to distribute the exponent thoroughly. sooner or later, I may post another question!!

OpenStudy (zzr0ck3r):

for sure:)